### v1

# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

Abstract
Permutation tests are a powerful and flexible approach to inference via resampling. As computational methods become more ubiquitous in the statistics curriculum, use of permutation tests has become more tractable. At the heart of the permutation approach is the exchangeability assumption, which determines the appropriate null sampling distribution. We explore the exchangeability assumption in the context of permutation tests for multiple linear regression models. Various permutation schemes for the multiple linear regression setting have been previously proposed and assessed in the literature. As has been demonstrated previously, in most settings, the choice of how to permute a multiple linear regression model does not materially change inferential conclusions. Regardless, we believe that (1) understanding exchangeability in the multiple linear regression setting and also (2) how it relates to the null hypothesis of interest is valuable. We also briefly explore model settings beyond multiple linear regression (e.g., settings where clustering or hierarchical relationships exist) as a motivation for the benefit and flexibility of permutation tests. We close with pedagogical recommendations for instructors who want to bring multiple linear regression permutation inference into their classroom as a way to deepen student understanding of resampling-based inference.
1 Introduction
Statistical inference is based on modeling the variability inherent in a dataset. Many of the models and analysis methods taught in a standard undergraduate statistics curriculum rely on asymptotic normal theory, with the theoretical underpinnings driven by the Central Limit Theorem. However, permutation tests are becoming increasingly popular because they provide a flexible approach for a wide scope of problems. Unlike methods base on the Central Limit Theorem, permutation tests do not generally require distributional or sample size assumptions. They do, however, require exchangeability, an idea which comprises much of the substance of the paper and which will be introduced in Section 1.1.
Permutation tests were among the first inferential tests conceived and used widely (Fisher 1935), in the context of categorical data. Permutation tests have since been expanded to cover many different modeling contexts and are often presented as a way to deepen an understanding of sampling distributions and normal theory methods (Ismay and Kim 2020; Chance and Rossman 2021; Baumer, Kaplan, and Horton 2024; Çentinkaya-Rundel and Hardin 2023). Normal inference with simple least squares regression requires that the data are independent, follow a linear model, and have error terms which are normal with equal variance. Permutation tests allow for inference on the least squares model where the normality condition can be dropped. The computational application of a permutation test in the simple linear model case requires the analyst to permute, equivalently, either the explanatory or the response variable before re-fitting the least squares model under the null hypothesis. However, the simple linear regression model is somewhat limited in the types of problems it can handle, and there has been a recent push to infuse the statistics curriculum with multivariate thinking (Carver et al. 2016).
Myriad approaches to permutation approaches in the multiple linear regression (MLR) model have been proposed in the literature. Manly (1997) and Winkler et al. (2014) summarize and synthesize a variety of MLR permutation methods. In N. Draper and Stoneman (1966), the treatment variable of interest is shuffled, while Manly (1986) shuffles the response variable itself. Permuting model residuals are considered in multiple papers. Freedman and Lane (1983) permute residuals from a null model, ter Braak (1992) permutes residuals from a full model, Kennedy (1995) permutes values from a model that residualizes both the response variable and the treatment variable, and Huh and Jhun (2001) permutes under conditions of a block structure experimental design. Still and White (1981) consider a special ANOVA case with interaction.
MLR provides a foundation for multivariate thinking that is both accessible to undergraduates and often taught in many introductory statistics classes and second courses in applied statistics. Permutation methods are ubiquitous in graduate studies, where student build up intuition, theory, and computation. Some aspects of permutation testing in the MLR setting are not fully straightforward. As an example, the “permute” step of a permutation test in the MLR is not as obvious as in the simple linear regression case, with the determination of how to permute depending on the underlying structure of the data, the specific hypotheses begin tested, and the assumption of exchangeability.
While not necessarily straightforward, working through some of the permutation options for MLR affords an understanding of exhangeability and permutation tests more generally that sets up students for understanding permutation tests in more complex settings. We review some common MLR permutation methods and motivate how and why exploration of permutation-based inference for multiple regression models is valuable.
1.1 Exchangeability
At the foundation of valid permutation tests is the condition of exchangeability (Pitman 1937; Good 2002). While there are typically no distributional or sample size restrictions on a permutation test, how to incorporate the exchangeability condition can sometimes be difficult (Welch 1990). D. Draper et al. (1993) gives a detailed account of the issues of exchangeability, including when more complicated settings where the data have serial correlation.
We use the following definition of exchangeability throughout:
Data are exchangeable under the null hypothesis if the joint distribution from which the data came is the same before permutation as after permutation when the null hypothesis is true.
Note, for example, if a confounding effect is present in a MLR model, the data are often not exchangeable, even when the null hypothesis is true. Consider an example whose goal is to model performance in college as the response variable ($Y$) based on whether the student took Advanced Placement (AP) courses in high school ($X_{2}$, the treatment variable) and the student’s family income level ($X_{1}$). It has been well documented that work on AP exams is associated with socio-economic status (Kolluri 2018). Assume that the null hypothesis is true; that is, having taken the AP exam is not linearly related to performance in college after controlling for family income.
In that setting, permuting the treatment variable (taking AP courses or not) creates a null sampling distribution with the unintended side-effect that the treatment variable ($X_{2}$) is no longer associated with income level ($X_{1}$). Because treatment and income are associated in the real data, the permuted data do not retain the observed dependency structure in the original data. That is, the relevant relationship is lost after permutation and the exchangeability condition is violated.
2 Permutation Tests in MLR
2.1 Notation
In order to work through the details of the permutation methods and corresponding exchangeability constraints, we provide notation and model specifications to describe the MLR model.
The population model of interest is
$E[Y]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2},$
where $X_{2}$ is the variable of interest in predicting $Y$, and $X_{1}$ is a nuisance variable. The $\cdot$ (dot) notation indicates the “other” variables included in the model.
The model estimated from the original dataset is given by
$\hat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
where $b_{1\cdot 2}$ is the sample coefficient on $X_{1}$ given that $X_{2}$ is in the model, and $b_{2\cdot 1}$ is the sample coefficient on $X_{2}$ given that $X_{1}$ is in the model.
A variable that is permuted is denoted by an asterisk. For example,
$Y^{*}\sim X_{1}+X_{2}$
means that permuted $Y$ values are regressed on the original $X_{1}$ and $X_{2}$. The corresponding model would be based on permuted values; therefore, the estimated coefficients would be indicated by asterisks: $b_{0\cdot 1,2}^{*}$, $b_{1\cdot 2}^{*}$, and $b_{2\cdot 2}^{*}.$
In Section 2.2.3, the permutation happens on residuals from a reduced model. That is, the model is first fit only on $X_{1}$, producing coefficients $b_{0\cdot 1}$ and $b_{1}$. Then residuals are calculated using
$R_{Y\cdot 1}=Y-(b_{0\cdot 1}+b_{1}X_{1}).$
The permuted residuals are added back to the single variable model, arriving at new $Y^{*}$ values which are then regressed on $X_{1}$ and $X_{2},$ the covariates.
$Y^{*}:=b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1}$
2.2 Permuting specifications
Unlike simple linear regression, with MLR there is no obvious choice of how to permute. In what follows, we discuss four different permutation methods and their exchangeability conditions. The goal of permuting is to create a null sampling distribution of the statistic of interest, here $b_{2\cdot 1},$ so thata we may infer the variability of the statistic. That is, a distribution of the statistics under the setting where the response $Y$ and the $X_{2}$ variable are not linearly related. But also, the null distribution needs to be created under the exchangeability condition—that the permuting only leads to a change in the relationship between $Y$ and $X_{2}$ and not to a change to any of the other variable relationships in the linear model. All four of the permuting schemes set the null hypothesis to $H_{0}\!\!:\beta_{2\cdot 1}=0.$ See Manly (1997) for an excellent summary of the different permutation methods within the MLR model.
Statistics that are independent of all unknown parameters are called pivot statistics. In our setting, we use a t-statistic (which is pivotal) as our test statistic for two reasons. First of all, it allows for comparison across all methods, even the method that permutes the residuals of the full model and does not have $\beta_{2\cdot 1}=0$ in the test statistic. Secondly, the advantages of pivot statistics are well-established (Hall and Wilson 1991; Westfall and Young 1993).
2.2.1 Permute $Y$
At first glance, it might seem like permuting the response variable would be a good way to break the relationship between $Y$ and $X_{2}$ (Manly 1986, 1997). Indeed, permuting $Y$ will break the relationship between $Y$ and $X_{2}$ which will force the null hypothesis to be true. However, permuting $Y$ will also break the relationship between $Y$ and $X_{1},$ which may not be acceptable if we need to preserve the relationship to mirror the original data structure. If we permute $Y,$ then exchangeability is met only if $Y$ and $X_{1}$ are uncorrelated in the original dataset.
Summary of permuting $Y$
Regression:
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $Y$. Note that when $Y$ is permuted, the original relationship between $X_{1}$ and $X_{2}$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $Y$, however, is problematic under general conditions where $X_{1}$ and $Y$ are not independent. It is worth pointing out that underlying all of the methods (including Ordinary Least Squares (OLS)), the observations are assumed to be independent of one another.
Table 1: Different permutation schemes, variable relationships that are broken, and variable relationships that are preserved. Note that all methods require the observations to be independent of one another.
Permutation Broken Relationships Preserved Relationships
Permute $Y$ $X_{1}$ $Y$ $X_{1}$ $X_{2}$
Manly (1986), Manly (1997) $X_{2}$ $Y$
Permute $X_{2}$ $X_{1}$ $X_{2}$ $X_{1}$ $Y$
N. Draper and Stoneman (1966) $X_{2}$ $Y$
Permute reduced model residuals $X_{2}$ $Y$ (if $X_{1}$ $X_{2}$ are not highly collinear) $X_{1}$ $X_{2}$
Freedman and Lane (1983) $X_{1}$ $Y$
Permute full model residuals None $X_{1}$ $X_{2}$
ter Braak (1990), ter Braak (1992) $X_{1}$ $Y$
$X_{2}$ $Y$
2.2.2 Permute $X_{2}$
In order to maintain the relationship between $Y$ and $X_{1}$, we might consider permuting $X_{2}$ instead of $Y$ (N. Draper and Stoneman 1966). Indeed, the permutation distribution created from permuting $X_{2}$ will force the null hypothesis to be true. However, permuting $X_{2}$ has the side effect that the relationship between $X_{1}$ and $X_{2}$ will be broken in the permuted data. If the data come from, for example, a randomized clinical trial (where $X_{2}$ is the treatment variable), then $X_{1}$ and $X_{2}$ will be independent in the original dataset, and permuting of $X_{2}$ will not violate the exchangeability condition. If $X_{1}$ and $X_{2}$ are correlated in the original dataset, as in our AP exam and socio-economic status example, then permuting $X_{2}$ violates the exchangeability condition.
Summary of permuting $X_{2}$
Regression:
$Y\sim X_{1}+X_{2}^{*}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $X_{2}$. Note that when $X_{2}$ is permuted, the original relationship between $X_{1}$ and $Y$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $X_{2}$, however, is problematic under general conditions.
2.2.3 Permute reduced model residuals
In order to address some of the concerns of the first two MLR permutation methods, Freedman and Lane (1983) propose a permutation scheme based on residuals. The permutation preserves the relationship between $X_{1}$ $X_{2}$ as well as the relationship between $X_{1}$ $Y$. However, in order for the relationship between $X_{2}$ $Y$ to be broken (i.e., to obtain a null sampling distribution for the test of $H_{0}:\beta_{2\cdot 1}=0),$ $X_{1}$ and $X_{2}$ must not be highly collinear.
Summary of permuting reduced model residuals
Regression:
$Y^{*}:=b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1}$
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that when permuting the reduced model residuals, the original relationships between both $X_{1}$ and $X_{2}$ and additionally between $X_{1}$ and $Y$ are preserved (which is what we want in terms of exchangeability). The null hypothesis is only true if the relationship between $X_{2}$ and $Y$ is broken, and that happens only when $X_{1}$ and $X_{2}$ are not highly collinear.
It is not immediately obvious that a high correlation between $X_{1}$ and $X_{2}$ leads to a non-null distribution. We provide evidence for the claim in two ways: by describing the permutation scheme and by giving a sketch for the proof of the claim.
Claim. The correlation between $Y^{*}$ and $X_{2}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
  1. (1)
    The permutation scheme is designed to create new values for the coefficient on $X_{2}$ under the setting that $X_{2}$ and $Y$ are uncorrelated. Repeated sets of the “permuted” $Y^{*}$ values are generated by adding noise (permuted residuals) to the fitted values from the model on $X_{1}$ only. At first glance, it seems as though the method creates $Y^{*}$ values that are associated with $X_{1}$ (because they built from the $X_{1}$-model fitted values) and not associated with $X_{2}$ (because they are built from a model that completely ignores $X_{2}$). However, if $X_{1}$ and $X_{2}$ are highly correlated, then $Y^{*}$ values that are correlated with $X_{1}$ will naturally be correlated with $X_{2}$.
  2. (2)
    In Appendix A we give a proof sketch outlining the derivation for the covariance between $Y^{*}$ and $X_{2}$. Although the proof is only outlined, the dependence on the correlation between $X_{1}$ and $X_{2}$ is clear. The approximate value of $\rho(Y^{*},X_{2})$ derived in Appendix A is validated empirically (results not shown); see Ye (2023) for empirical results.
2.2.4 Permute full model residuals
As an extension to Freedman and Lane (1983), ter Braak’s (1990, 1992) permutation method permutes the residuals from the full model. The method has a bootstrap flavor, but the permutation of the residuals is done without replacement (whereas bootstrapping is done with replacement). The underlying technical condition for ter Braak’s method is that:
$F\bigg{(}\frac{b_{2\cdot 1}^{*}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}\bigg{)}\approx F\bigg{(}\frac{b_{2\cdot 1}-\beta_{2\cdot 1}}{SE(b_{2\cdot 1})}\bigg{)},$ (1)
where $F$ is the cumulative distribution function of the random variable described in the argument of the function.
The sampling distribution of $t^{*}=\frac{b^{*}_{2\cdot 1}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}$ approximates the sampling distribution of $t=\frac{b_{2\cdot 1}-\beta_{2\cdot 1}}{SE(b_{2\cdot 1})}$ due to the theoretical underpinnings from bootstrapping. That is, the variability of $b_{2\cdot 1}^{*}$ around $b_{2\cdot 1}$ mimics the variability of $b_{2\cdot 1}$ around $\beta_{2\cdot 1}.$
Under Equation (1), we can use the $t^{*}$ distribution constructed from many permutations of the same dataset to carry out a hypothesis test. If $H_{0}\!\!:\beta_{2\cdot 1}=0$ is true, then $t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$ would be a likely value in the $t^{*}$ distribution, corresponding to a non-significant p-value, resulting in a failure to reject $H_{0}.$
However, if $H_{A}\!\!:\beta_{2\cdot 1}\neq 0$ is true and, say, $\beta_{2\cdot 1}=47,$ then we would expect $\hat{t}=\frac{b_{2\cdot 1}-47}{SE(b_{2\cdot 1})}$ to lie well within the $t^{*}$ distribution, while $t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$ would lie on the margins, leading to a small p-value that concludes the test by rejecting $H_{0}.$
Furthermore, ter Braak (1990) synthesizes equations from various authors – Efron (1982) for the bootstrap, along with Cox and Hinkley (1974) and Lehmann and D’Abrera (1975) for the permutation – to make the following statements regarding the expected value and variance of the estimated slope coefficients in the bootstrapping versus permutation settings. $b_{2\cdot 1}^{+}$ corresponds to the coefficient derived from a standard bootstrap (full model), while $b_{2\cdot 1}^{*}$ corresponds to the permutation of the residuals under the full model.
$E^{+}(b_{2\cdot 1}^{+})$ $=E^{*}(b_{2\cdot 1}^{*})=b_{2\cdot 1}$ (2)
$var^{+}(b_{2\cdot 1}^{+})$ $=(1-1/n)var^{*}(b_{2\cdot 1}^{*}).$ (3)
Equation (2) indicates that the expected values of $b_{2\cdot 1}^{+}$ and $b_{2\cdot 1}^{*}$ are both $b_{2\cdot 1}$. Equation (3) suggests that the variance of $b_{2\cdot 1}^{+},$ the bootstrapped estimate, is smaller than the variance of $b_{2\cdot 1}^{*}.$ Hence, ter Braak uses Equations (2) and (3), along with the order property that $b_{2\cdot 1}^{+}$ and $b_{2\cdot 1}^{*}$ differ by $O(1/n)$ in second or higher order moments, to justify his proposal of the full model residual permutation strategy.
Summary of permuting full model residuals
Regression:
$Y^{*}:=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}+R^{*}_{Y\cdot 1,2}$
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b^{*}_{2\cdot 1}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that permuting the residuals under the full model allows all of the exchangeability conditions to hold. The new model does not force the null hypothesis to be true which is why the test statistic measures the deviation of the permuted coefficients (on $X_{2}$) to the original data model coefficient (on $X_{2}$) instead of comparing to zero.
2.3 Why permutation test?
It turns out that even when both the normality conditions and the exchangeability are violated, all four of the aforementioned permutation test methods do reasonably well at controlling Type I errors (Anderson and Legendre 1999; Winkler et al. 2014). Only when the error terms come from extremely skewed distributions do the permutation tests demonstrate substantially higher power than the normal theory methods (Anderson and Legendre 1999). There is strong evidence that in order to violate alpha-level testing, the data must violate the technical conditions to an extreme degree (Anderson and Legendre 1999; Tantawanich 2006; Winkler et al. 2014). Our own simulations (not shown) give the same results that despite differences across exchangeability conditions, the different permutation schemes presented above do not result in substantially different inferential outcomes.
If we anticipate that results will be indistinguishable, why is an understanding of permutation methods (and exchangeability) important?
  1. 1.
    We believe that exchangeability is a valuable concept for students to understand, as it undergirds key foundational knowledge of statistical inference.
  2. 2.
    There do exist settings where mathematical and permutation approaches produce different inferential results, even within the linear model framework (e.g., see Section 4). Knowledge of permutation-based approaches (plus computational skills) allows inference in such areas where linear model assumptions don’t hold, and permutation tests may be straightforward to implement in situations where typical parametric tests may exist but be extremely complicated. As long as permuted draws can be made from the appropriate null world, permutation tests can be used to make inferences without parametric assumptions.
3 Full example analysis
A full simulation study comparing permutation methods is both redundant (of the literature, see Anderson and Legendre (1999) and Winkler et al. (2014)) and outside of the goals of our paper. However, we provide an analysis of a single real dataset to demonstrate the methods and compare the resulting p-values.
The data come from the National Bridge Inventory (NBI) [^1] https://www.fhwa.dot.gov/bridge/nbi.cfm. Subject to the National Bridge Inspection Standards, each state is responsible for preparing and maintaining an inventory of all bridges. NBI provides data for New York in 2023222https://www.fhwa.dot.gov/bridge/nbi/2023/delimited/NY23.txt with a complete coding guide [^3] https://www.fhwa.dot.gov/bridge/mtguide.pdf. In what follows, we examine a small random sample of bridges in New York that go over highways and are made of either steel or concrete. Our interest is in modeling improvement cost (proposed cost of improvement) as a linear function of average daily traffic volume ($X_{1}$) and the material ($X_{2}$) of the bridge (either concrete or steel). Figure 1 provides a visual representation of the linear relationship between cost and volume, with a shift due to the type of material.
In order to assess whether or not the type of material (steel or concrete) is a significant predictor variable in the multiple linear regression model of improvement cost on volume and type of material, we run five different hypothesis tests (all with the same null hypothesis that the coefficient on type of material is zero, i.e., $H_{0}\!\!:\beta_{2\cdot 1}=0$).
$E[\texttt{cost}]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}\texttt{volume}+\beta_{2\cdot 1}\texttt{material}$
The five models are OLS and the four permutation methods described above in Section 2. Table 2 displays the p-values calculated by each of the five different hypothesis tests. Some important aspects to note from the analysis are:
  • The average daily traffic volume is significant. (The p-value is not shown, but the relationship is seen in Figure 1).
  • The two explanatory variables (average daily traffic volume and type of material) are potentially associated, as seen in Figure 2.
  • The data are not symmetrically distributed about the regression lines, violating the normality condition of ordinary least squares.
  • All of the models require that the observations are independent.
Despite multiple potentially important model condition violations (e.g., normality of residuals and homoscedasticity of residuals), all five p-values (see Table 2) are on the same order of magnitude and produce the same inferential conclusions about the coefficient on type of material ($X_{2}$).
Figure 1: For a small sample of bridges in New York state, the cost for improvement seems to be a function of both the average daily traffic volume of the bridge as well as the bridge material (steel or concrete).
Table 2: A comparison of the resulting p-values obtained from applying ordinary least squares and each of the permutation methods to data on bridges in New York state with 2000 permutations. The inference seeks to understand the effect of construction material (steel or concrete) on cost of repairs in a regression model adjusted by traffic volume.
permutation type p-value
treatment variable (Draper & Stoneman) 0.14
response variable (Manly) 0.16
reduced model residuals (Freedman & Lane) 0.13
full model residuals (ter Braak) 0.13
OLS 0.15
Figure 2: The average daily traffic volume seems potentially associated with the type of material used in the bridge, with concrete bridges having a higher median traffic volume.
4 Extensions to account for clustered data
The bridges example demonstrates how permutation testing can be used to undertake inference in standard multiple regression model settings. As we have seen above, ideas of exchangeability are reasonably straightforward in the MLR model, and we can walk through the details to understand how the exchangeability conditions impact the appropriate conclusions (see Table LABEL:tbl-perms).
However, permutation methods are particularly attractive since they can be adapted to more complicated settings where exchangeability might not be as straightforward. In particular, one of the conditions for inference in multiple regression is that the observations (or more accurately, the residuals) are independent of one other. For the bridges setting, we might speculate that the observations are not independent. For example, consider bridges in a particular area (e.g., a city) that reflect underlying features, which might systematically affect traffic flow and/or cost (e.g., availability of resources or contracts with particular construction companies). We might imagine that two bridges in the same city would share a common construction market and patterns of use that are associated with the cost.
Violations of independence show up in many types of studies. For example, consider a situation where, for convenience or design purposes, families might be sampled rather than individuals. Principled and flexible methods have been developed for multiple linear regression models in the setting where the observations are clustered (see for example Laird and Ware (1982)). Random intercept models, where each observational unit (e.g., family) is assumed to have an underlying (unobserved) random level, are commonly used to account for clustering. However, such mixed effects models are both complicated (mixed effects models are not typically seen in undergraduate curricula) and may require sufficient sample size to assume asymptotic normality of regression coefficients (Maas and Hox 2005).
4.1 Cluster permuting
We consider an example where we are interested in predicting an outcome ($Y$) based on a quantitative predictor ($X_{1}$) and a dichotomous treatment ($X_{2}$) where some of the families have two individuals within the sample. An adaption of the permutation tests that accounts for clustering within observational unit is a computational alternative to the mathematical approach, which maintains appropriate Type I error rate (Winkler et al. 2014, 2015). Here we explore the clustered permutation idea with three different data scenarios, representing independent observations and two different types of clustering that violate the independent observation requirement.
4.1.1 Permuting independent observations
Figure 3 displays a simple example where the observations are independent and there is no clustering structure. There are eleven subjects, each from a different family; six of the subjects are exposed and five of the subjects are control samples.
In the example, we permute $X_{2}$, akin to the method of N. Draper and Stoneman (1966). We note that because we are permuting, there are always five exposed observations and six control observations, leading to (${11\choose 6}$ = 462 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a multiple linear regression on the quantitative predictor and the permuted treatment and use the resulting null distribution of the t-statistics to assess the significance of the treatment variable.
Figure 3: Each color and row represents a family structure. There are 11 individuals, none of them are in the same family. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure.
4.1.2 Permuting homogeneous clustered observations
In the homogeneous clustered (e.g., same value of the treatment within a family) data scenario, there are multiple observations of the same type within the observational unit (see Figure 4). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With homogeneous clustered observations, the permutation scheme must align with the family structure. Here, each family has only one treatment (either exposure or control), so we permute the treatments with the constraint that three of the single-person families and one of the two-person families are control. For the homogeneous clustered example and scenario, there are (${5\choose 3}\cdot{3\choose 1}$ = 30 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). After each permutation, we fit a multiple linear regression on the quantitative predictor and the permuted treatment. The resulting null distribution of the t-statistics is used to assess the significance of the treatment variable.
Figure 4: Each color and row represents a family structure. There are 5 individual “families” and three families with two members each. In the original structure treatment happens by family (e.g., air pollution near / not near the family home). Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. The families are permuted together so as to keep intact the original scenario of treatment within family.
4.1.3 Permuting heterogeneous clustered observations
In the heterogeneous clustered (e.g., each family has members with both values of the dichotomous treatment) data scenario, families with two observations have individuals of each type of treatment observed within each observational unit (see Figure 5). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With heterogeneous clustered observations, the permutation scheme must align with the family structure. Here, the treatment on individual families will be permuted as if the observations were independent (i.e., a permutation with the constraint that two individuals are exposed and three are control). The permutation of observations in the two-member families happens within the family. That is, with equal probability, the treatment levels are either kept the same or swapped, within each family.
For the heterogeneous clustered scenario, there are (${5\choose 3}\cdot 2^{3}$ = 80 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). After each permutation, we fit a multiple linear regression on the quantitative predictor and the permuted treatment. The resulting null distribution of the t-statistics is used to assess the significance of the treatment variable.
Figure 5: Each color and row represents a family structure. There are 5 individual “families” and three families with two members each. In the original structure, for the families of size two, exactly one member of the family is exposed, and one member of the family is in the control group. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. For the families of size two, the permutation happens within the family so as to keep intact the original structure of both treatments across the family.
4.2 Cluster simulations
We undertook a limited series of simulation studies to explore the behavior of permutation tests when the design structure determined a violation of the independence condition. These simulations exploited the fact that ignoring clustering for the homogeneous clusters yields standard errors that are too small and that doing so for the heterogeneous cluster design, where family members serve as their own control, yields standard errors that are too large (Cannon et al. 2001).
The heart of our simulations is well described by Winkler et al. (2015) who focuses on
any dataset with known dependence among observations. In such cases, some permutations, if performed, would create data that would not possess the original dependence structure, and thus, should not be used to construct the reference (null) distribution. To allow permutation inference in such cases, we test the null hypothesis using only a subset of all otherwise possible permutations, i.e., using only the rearrangements of the data that respect exchangeability, thus retaining the original joint distribution unaltered.
For each model/scenario combination, 2000 simulations were generated with n = 20 subjects per group (more subjects than in Figure 3, Figure 4, and Figure 5). For the clustering scenarios, half of the subjects were singletons and the others shared a family member within the group (homogeneous) or across groups (heterogeneous). Within each simulation the null distribution was created using 2000 permutations.
Table 3: Type I error rates for different scenarios (2000 simulations, each with 2000 permutations for each simulation). Models include multiple linear regression (lm), multiple linear mixed effects (lme), permutation not accounting for clustering (naïve), and permutation accounting for clustering (correct). Forms of clustering are independent (yellow), homogeneous (white), and heterogeneous (blue). The dichotomous treatment variable was permuted (Draper and Stoneman, 1996).
Model/Scenario Type I error rate (95% CI) Results
lm independent 0.052 (0.043-0.063) matches desired alpha level
correct permutation independent 0.054 (0.045-0.065) matches desired alpha level
lm homogeneous 0.08 (0.068-0.092) anticonservative
lme homogeneous 0.048 (0.04-0.059) matches desired alpha level
naïve permutation homogeneous 0.081 (0.069-0.094) anticonservative
correct permutation homogeneous 0.055 (0.045-0.066) matches desired alpha level
lm heterogeneous 0.025 (0.019-0.033) conservative
lme heterogeneous 0.044 (0.036-0.054) matches desired alpha level
naïve permutation heterogeneous 0.03 (0.023-0.038) conservative
correct permutation heterogeneous 0.054 (0.045-0.065) matches desired alpha level
Table LABEL:tbl-simresults displays the results from the simulation study. As expected, the models that do not appropriately account for clustering are anticonservative (“lm homogeneous”, “naïve permutation homogeneous”) or overly conservative (“lm heterogeneous” and “naïve permutation heterogeneous”); the models are given in boldface. Methods that appropriately account for the clustering (e.g., “lme” or “correct permutation”) yielded the desired Type I error rate. When there are cluster dependencies, the naïve method is not advised.
5 Discussion and pedagogical recommendations
Our work presents some of the existing literature on permutation tests in the multiple linear regression setting. While it is not immediately obvious how to best permute an MLR model, it turns out that, generally, the different methods perform similarly with respect to Type I errors (and power). However, the difficult and extremely powerful concept of exchangeability can be accessed through the MLR setting in the classroom. After digging into the exchangeability ideas for the MLR case, we have found that students are able to apply a permutation approach to more complicated data settings (which are often encountered in real applications), like clustered observations. This guidance is consistent with the advice of Chance et al. (2024), who hoped “to help instructors understand the differences in the simulation strategies to better inform their own decisions of how to adapt a simulation approach for their classes and to better respond to student questions that may arise”.
Another compelling rationale to devote time to these questions is the concept that “the test follows from the design”:
There is a clear logical link between the statistical test we use and the experimental design we opted for: using a permutation test is entirely warranted by the random assignment of participants to two equal-sized groups. Stressing the link between experimental design and statistical inference – rather than considering them separately – is of huge pedagogical, as well as practical, use, I believe (Vanhove 2015).
We close with some practical guidance about how to bring exchangeability and permutation tests into the classroom:
  • Teach permutation tests for MLR. Talk to your students about the choice they will have to make (permute $Y$? $X_{2}$? residuals?) and that all modeling contains choices. Choices are not usually objective or unbiased, so an ability to defend their choice is what gives them power as a statistician.
  • Create a simulation program (Rmd/qmd file or Shiny App) where students can change the error structure and discover the desired Type I error rates are not achieved only when the errors are particularly egregious.
  • Describe the difference between creating a null sampling distribution (i.e., making sure that the null hypothesis is true) and establishing that exchangeability holds (i.e., making sure that the Type I error rates will be accurate).
  • Present some or all of the MLR methods described here (with more in Winkler et al. (2014)) and have students fill out a blank Table LABEL:tbl-perms.
  • Ask your students what it means that “the test follows from the design.” Use the cluster scenarios and apply the opposite permutation scheme. Have them compare the permutation scheme and the experimental design to figure out what makes most sense in terms of the relationship of the permuting and the design.
  • Come up with scenarios (or use our cluster scenarios!) where the students can figure out the correct permutation schemes. See an example from JH’s class using a stratified two-sample test. https://st47s.com/Math154/Notes/permschp.html#macnell-teaching-evaluations-stratified-two-sample-t-test
  • Find permutation tests in the literature and have students describe how the permutation scheme upholds (or doesn’t!) the exchangeability condition.
  • Find theoretical work that proves exchangeability and demonstrate the important mapping of the theory to the applied problems (or maybe the exchangeability is not vital to the performance of the method?).
The ideas we have put forward can be taught at a variety of levels. Whether to permute $Y$ or $X_{2}$ can be taught as early as introductory statistics where students have done some hypothesis testing and some multivariate modeling. Graduate students can think carefully about how to prove that conditions of exchangeability are met. But the group of students who might most benefit from the ideas we’ve presented include upper level undergraduates and early graduate students who are focused on understanding which models are best in which settings and how to differentiate those settings. We believe that the discussions which will come from our presentation will be interesting and fruitful for broad statistical learning.
6 Acknowledgements
This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. We thank Shiya Cao and Lindsay Poirier for pointing us toward the bridges data. We thank Shiya Cao for useful comments on an earlier draft of the manuscript.
7 Disclosure statement
The authors report there are no competing interests to declare.
8 Reproducibility statement
The code used to undertake the Bridges permutation analysis in Section 3 has been made available at https://osf.io/7zcfu/.
Johanna Hardin: https://orcid.org/$0$-0001-6251-1955
Nicholas Horton: https://orcid.org/$0$-0003-3332-4311
A Appendix
Claim. The correlation between $Y^{*}$ and $X_{2}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
Proof sketch.
Throughout the proof sketch, there are places where we have simplified the argument by considering statistics to be fixed parameters (not an unreasonable approximation under large sample sizes where Slutsky’s Theorem holds (Slutsky 1925)).
$\rho(Y^{*},X_{2})$ $=\frac{cov(Y^{*},X_{2})}{\sqrt{var(Y^{*})var(X_{2})}}$
Breaking down each part of the correlation between $Y^{*}$ and $X_{2}$, we compute both $cov(Y^{*},X_{2})$ and $var(Y^{*}).$
$cov(Y^{*},X_{2})$ $=cov(b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1},X_{2})$
$=cov(b_{0\cdot 1},X_{2})+cov(b_{1}X_{1},X_{2})+cov(R^{*}_{Y|1},X_{2})$
$\approx cov(b_{1}X_{1},X_{2})$ (4)
$\approx b_{1}\cdot cov(X_{1},X_{2})$ (5)
$=b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}$
Note that $b_{0}\cdot 1$ and $b_{1}$ are a random variables because they are a statistics, so Equations (5,6,7) are all approximate. In Equation (4), we assume that $b_{0\cdot 1}$ and $R^{*}_{Y\cdot 1}$ are both independent of $X_{2}$, which leads to $cov(b_{0\cdot 1},X_{2})=cov(R^{*}_{Y\cdot 1},X_{2})=0$.
$var(Y^{*})$ $=var(b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y|1})$
$\approx(b_{1})^{2}var(X_{1})+var(R^{*}_{Y|1})$ (6)
$var(R^{*}_{Y\cdot 1})$ is broken down into pieces.
$var(R^{*}_{Y\cdot 1})$ $=var(R_{Y\cdot 1})$
$=var(Y-b_{0\cdot 1}-b_{1}X_{1})$
$\approx var(Y)+(b_{1})^{2}var(X_{1})-2b_{1}cov(Y,X_{1})$ (7)
$cov(Y,X_{1})$ $=cov(\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2}+\varepsilon,X_{1})$
[…]
$=\beta_{1\cdot 2}var(X_{1})+\beta_{2\cdot 1}cov(X_{2},X_{1})$ (8)
Plugging $cov(Y,X_{1})$ into $var(R^{*}_{Y\cdot 1})$ and $var(R^{*}_{Y\cdot 1})$ into $var(Y^{*})$, we can approximate $var(Y^{*}).$
[…]
$=2b_{1}(b_{1}-\beta_{1\cdot 2})var(X_{1})+var(Y)-2b_{1}\beta_{2\cdot 1}cov(X_{2},X_{1})$
$=2b_{1}[(b_{1}-\beta_{1\cdot 2})var(X_{1})-\beta_{2\cdot 1}cov(X_{1},X_{2})]+var(Y)$
Putting it all together gives:
$\rho(Y^{*},X_{2})$ $=\frac{cov(Y^{*},X_{2})}{\sqrt{var(Y^{*})var(X_{2})}}$
$\approx\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}}{\sqrt{var(Y^{*})var(X_{2})}}$
$=\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})}}{\sqrt{var(Y^{*})}}$
$=\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})}}{\sqrt{2b_{1}[(b_{1}-\beta_{1\cdot 2})var(X_{1})-\beta_{2\cdot 1}cov(X_{1},X_{2})]+var(Y)}}.$ (9)