### v2

# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

Abstract
Permutation tests are a powerful and flexible approach to inference via resampling. As computational methods become more ubiquitous in the statistics curriculum, use of permutation tests has become more tractable. At the heart of the permutation approach is the exchangeability assumption, which determines the appropriate null sampling distribution. We explore the exchangeability assumption in the context of permutation tests for multiple linear regression models, including settings where the assumption is not tenable. Various permutation schemes for the multiple linear regression setting have been proposed and assessed in the literature. As has been demonstrated previously, in most settings, the choice of how to permute a multiple linear regression model does not materially change inferential conclusions with respect to Type I errors. However, some violations (e.g., when clustering is not appropriately accounted for) lead to issues with Type I error rates. Regardless, we believe that understanding (1) exchangeability in the multiple linear regression setting and also (2) how it relates to the null hypothesis of interest is valuable. We close with pedagogical recommendations for instructors who want to bring multiple linear regression permutation inference into their classroom as a way to deepen student understanding of resampling-based inference.
1 Introduction
Statistical inference is based on modeling the variability inherent in a dataset. Many of the models and analysis methods taught in a standard undergraduate statistics curriculum rely on asymptotic normal theory, with the theoretical underpinnings driven by the Central Limit Theorem. However, permutation tests are becoming increasingly popular because they provide a flexible approach for a wide scope of problems. Unlike methods based on the Central Limit Theorem, permutation tests do not generally require distributional or sample size assumptions. They do, however, require exchangeability, an idea which comprises much of the substance of our paper and which will be introduced in Section 1.1.
Permutation tests were among the first inferential tests conceived and used widely (Fisher 1935) in the context of categorical data. Permutation tests have since been expanded to cover many different modeling contexts and are often presented as a way to deepen an understanding of sampling distributions and normal theory methods. Permutation methods have long been pervasive in the statistics curriculum for graduate studies, where students build up intuition, theory, and computation. However, the undergraduate statistics curriculum has historically not had space for extended explication of permutation methods, although many modern introductory textbooks do introduce the basics of permutation tests (Ismay and Kim 2020; Chance and Rossman 2021; Baumer, Kaplan, and Horton 2024; Çentinkaya-Rundel and Hardin 2023). Regression, on the other hand, plays a central role within undergraduate minor and major programs (American Statistical Association 2014).
Normal inference with simple least squares regression, also known as simple linear regression (SLR), requires that the data are independent, follow a linear model, and have error terms which are approximately normal with equal variance. Permutation tests are attractive, since they allow for inference on the least squares model where the normality condition does not hold (particularly with small sample sizes, since it may not be tenable to invoke the Central Limit Theorem).
The computational application of a permutation test in the SLR model case requires the analyst to permute, equivalently, either the predictor or the outcome variable before re-fitting the least squares model. (A third option, permuting the residuals, is discussed later in Section 3.3.3 and Section 3.3.4.)
However, the SLR model is somewhat limited in the types of problems it can handle, and there has been a recent push to infuse the statistics curriculum with multivariate thinking (Carver et al. 2016). Multiple linear regression (MLR) provides a foundation for multivariate thinking that is both accessible to undergraduates and often taught in many introductory statistics classes and second courses in applied statistics (see, for example, a discussion of the undergraduate statistics curriculum in which teaching exchangeability is explicitly referenced in Kennedy-Shaffer (2024)).
Unlike SLR, in the MLR model, what to permute is not immediately obvious within the inferential process. The “permute” step of a permutation test in the MLR model is not as obvious as in the SLR case, with the determination of how to permute depending on the underlying structure of the data, the specific hypotheses begin tested, and the assumption of exchangeability.
The statistics literature includes myriad proposals for different permutation approaches in the MLR model. Manly (1997) and Winkler et al. (2014) summarize and synthesize a variety of MLR permutation methods. In N. Draper and Stoneman (1966), the treatment variable of interest is shuffled, while Manly (1986) shuffles the outcome variable itself. Multiple authors consider permuting model residuals: Freedman and Lane (1983) permute residuals from a null model, ter Braak (1992) permutes residuals from a full model, Kennedy (1995) permutes values from a model that residualizes both the outcome variable and the treatment variable, and Huh and Jhun (2001) permute under conditions of a block structure experimental design. Still and White (1981) consider a special ANOVA case with interaction. Recent work has described a variety of robust permutation approaches (DiCiccio and Romano 2017; Helwig 2019a, 2019b).
While not necessarily straightforward, working through some of the permutation options for MLR provides an understanding of exchangeability, and permutation tests more generally, that sets up students for understanding permutation tests in more complex settings. Our goal is to present MLR permutation methods, explore additional complications, and motivate how and why an investigation into permutation-based inference for multiple regression models is valuable as a way to expand understanding of inference and statistical foundations.
1.1 Exchangeability
At the foundation of valid permutation tests is the condition of exchangeability (Pitman 1937; Good 2002). While there are typically no distributional or sample size restrictions on a permutation test, recognizing how to incorporate the exchangeability condition can sometimes be difficult (Welch 1990). D. Draper et al. (1993) gives a detailed account of the issues of exchangeability, including more complicated settings, such as serial correlation.
We use the following definition of exchangeability throughout:
Data are exchangeable under the null hypothesis if the joint distribution from which the data came is the same before a permutation as after a permutation when the null hypothesis is true.
For linear regression models with a single predictor (SLR), equivalent results will ensue no matter whether the outcome variable ($Y$) or the predictor variable ($X$) is permuted.
Carrying out a permutation procedure involves calculation of a statistic for each of the permuted samples, then comparing the observed statistic to the permuted distribution where the association between $Y$ and $X$ has been broken. Here it doesn’t matter whether the outcome $Y$ or the predictor $X$ is permuted: the results will be equivalent.
A typical choice of statistic is a t-statistic, where the estimated regression coefficient is divided by the estimated standard error (Janssen 1997; Konietschke and Pauly 2012). Such a pivotal statistic has attractive properties, including providing robustness to modest deviations from exchangeability.
Undertaking a permutation test is more complicated when a second predictor is of interest, thereby making the MLR the desired model.
1.2 Motivating example
Consider an example whose goal is to model performance in college as the outcome variable ($Y$) based on whether the student took Advanced Placement (AP) courses in high school ($X_{1}$, the “treatment” variable) and the student’s family income level ($X_{2}$, a potential confounder). It has been well-documented that work in AP courses is positively associated with socio-economic status (Kolluri 2018). Assume that the null hypothesis ($\beta_{1}=0$) is true; that is, having taken an AP course is not linearly related to performance in college after controlling for family income. For the moment, we assume that the families are independent and only one student is included from each family.
Understanding exchangeability is sometimes best understood when it is violated. Imagine now carrying out a test from the given example setting by permuting $Y$ for this MLR model. The permutation process would break the relationship between $X_{1}$ and $Y$ (as desired), would preserve the relationship between $X_{1}$ and $X_{2}$ (as desired), but would also unfortunately break the relationship between $Y$ and $X_{2}$.
Breaking the relationship between $Y$ and $X_{2}$ wouldn’t be an issue if $Y$ and $X_{2}$ were not associated. But if $X_{2}$ is a potential confounder of the relationship, then $Y$ and $X_{2}$ may be associated. The permutation procedure where the outcome $Y$ is permuted, therefore, is not exchangeable, even when the null hypothesis of interest is true (i.e., $\beta_{1}=0$). In this setting, permuting the outcome creates a permutation sampling distribution with the unintended consequence that the performance in college ($Y$) is no longer associated with income level ($X_{2}$). Because treatment and income are associated in the real world, the data with permuted $Y$ values do not retain the observed dependency structure that is in the original data. That is, a relevant relationship is lost after a permutation, which ends up violating the exchangeability condition.
Similar issues arise if the predictor $X_{1}$ is permuted (which breaks the relationship between $Y$ and $X_{1}$, as well as between $X_{1}$ and $X_{2}$, but maintains the relationship between $Y$ and $X_{2}$). In either case, more relationships are broken than intended.
What impact does violation of exchangeability have on our inferences? We might consider studies of Type I error rate (when the null hypothesis is true) as well as Type II error rate (when the alternative hypothesis is true). Prior research (Anderson and Legendre 1999; Winkler et al. 2014) has shown that hypothesis testing in MLR is quite robust to the choice of permutation, and, except in cases of extreme error distributions, meeting (or not) the exchangeability conditions (such as breaking the association between $Y$ and $X_{2}$) does not substantially impact the size or power of the test.
However, robustness to exchangeability violations does not hold for all types of exchangeability violations. Consider another example where the underlying data are clustered in some fashion. For example, consider a situation where, for convenience or design purposes, families might be sampled rather than individuals. So instead of having just one individual from each family, we have two students for some families.
One of the conditions for inference in multiple regression is that the observations (or more accurately, the residuals) are independent of one other. In the clustering setting, it may no longer be tenable to assume that individuals within a family are independent of others within the same family.
Principled and flexible methods have been developed for MLR models in the setting where the observations are clustered (see for example Laird and Ware (1982)). Random intercept models, where each observational unit (e.g., family) is assumed to have an underlying (unobserved) random level, are commonly used to account for clustering. However, such mixed effects models are both complicated (mixed effects models are not typically seen in undergraduate curricula) and may require sufficient sample size to assume asymptotic normality of regression coefficients (Maas and Hox 2005).
Permutation methods are particularly desirable with clustered data since they can be adapted to more complicated settings where exchangeability might not be as straightforward as it is in the MLR setting (D. Draper et al. 1993). In the next section we detail how violations of exchangeability in the cluster setting can have a substantial impact on inferential results.
2 When exchangeability affects performance
We revisit our motivating example where we are interested in predicting college performace as the outcome ($Y$) based on a quantitative predictor (family income, $X_{2}$) and a dichotomous treatment (AP course taking, $X_{1}$) where some of the families have two individuals within the sample. An adaption of the permutation tests that accounts for clustering within observational unit is a computational alternative to the mathematical (i.e., based on the Central Limit Theorem) approach, which maintains an appropriate Type I error rate (Winkler et al. 2014, 2015). Here we explore the clustered permutation idea with three different data scenarios, representing independent observations and two different types of clustering that violate the independent observation condition.
2.1 Cluster scenarios
2.1.1 Permuting independent observations
Figure 1 displays an example where the observations are independent and no clustering structure exists. There are eleven subjects, each from a different family; six of the subjects are exposed, $X_{1}$, (e.g., the exposure group takes an AP course) and five of the subjects are control samples (e.g., no AP course). A second variable, $X_{2}$, (e.g., income) is collected on each subject and the outcome variable, $Y$, (e.g., college performance) is measured.
In the example, we permute $X_{1}$, akin to the method of N. Draper and Stoneman (1966). We note that because we are permuting, there are always five exposed observations and six control observations, leading to (${11\choose 6}$ = 462 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a MLR for each permutation use the resulting null distribution of the t-statistics (i.e., the standardized slope coefficients for $X_{1}$) to assess the significance of the treatment variable.
Figure 1: Each color and row represents a family structure. There are eleven individuals, none of whom are in the same family. Note that in the permutation setting, the exposure group and the control group both retain the same number of individuals as in the original family structure.
2.1.2 Permuting homogeneous clustered observations
In the homogeneous clustered (i.e., all individuals within a family have the same value of the treatment variable) data scenario, there are multiple observations of the same type within the observational unit (see Figure 2). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With homogeneous clustered observations, the permutation scheme must align with the family structure. Here, each family has only one treatment (either exposure or control), so we permute the treatments with the constraint that three of the single-person families and one of the two-person families are control. For the homogeneous clustered example, there are (${5\choose 3}\cdot{3\choose 1}$ = 30 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a MLR for each permutation use the resulting null distribution of the t-statistics (i.e., the standardized slope coefficients for $X_{1}$) to assess the significance of the treatment variable.
Figure 2: Each color and row represents a family structure. There are five single-person “families” and three families with two members each. In the original structure, treatment happens by family. Note that in the permutation setting, the exposure group and the control group both retain the same number of individuals as in the original family structure. The families are permuted together so as to keep intact the original scenario of treatment within family.
2.1.3 Permuting heterogeneous clustered observations
In the heterogeneous clustered (e.g., each family has members with both values of the dichotomous treatment) data scenario, families with two observations have individuals of each type of treatment observed within each observational unit (see Figure 3). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With heterogeneous clustered observations, the permutation scheme must perserve the family structure. Here, the treatment on single-person families will be permuted as if the observations were independent (i.e., a permutation with the constraint that two individuals are exposed and three are control). The permutation of observations in the two-person families happens within the family. That is, with equal probability, the treatment levels are either kept the same or swapped, within each family.
For the heterogeneous clustered scenario, there are (${5\choose 3}\cdot 2^{3}$ = 80 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). After each permutation, we fit a multiple linear regression on the quantitative predictor and the permuted treatment. The resulting null distribution of the t-statistics is used to assess the significance of the treatment variable.
Figure 3: Each color and row represents a family structure. There are five single-person “families” and three families with two members each. In the original structure, for the families of size two, exactly one member of each family is exposed, and the other member of the family is in the control group. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. For the families of size two, the permutation happens within the family so as to keep intact the original structure of both treatments across the family.
2.2 Cluster simulations
Recall that in the independent MLR setting, even permutation methods that violate exchangeability did not show substantial performance declines (Anderson and Legendre 1999; Winkler et al. 2014). With the clustered family structure, we have constructed a completely different type of exchangeability violation due to the independence (or lack thereof) of observations. The independence exchangeability violation does matter (in terms of performance) as seen in what follows.
We undertook a limited series of simulation studies to explore the behavior of permutation tests when the design structure was intentionally specified to violate the independence condition. The simulations exploited the fact that ignoring clustering for the homogeneous clusters yields standard errors that are too small and that doing so for the heterogeneous cluster design, where family members serve as their own control, yields standard errors that are too large (Cannon et al. 2001).
The heart of our simulations is well described by Winkler et al. (2015), who focus on
any dataset with known dependence among observations. In such cases, some permutations, if performed, would create data that would not possess the original dependence structure, and thus, should not be used to construct the reference (null) distribution. To allow permutation inference in such cases, we test the null hypothesis using only a subset of all otherwise possible permutations, i.e., using only the rearrangements of the data that respect exchangeability, thus retaining the original joint distribution unaltered.
For each model/scenario combination, 2000 simulations were generated with $n=20$ subjects per treatment group (more subjects than in Figure 1, Figure 2, and Figure 3). For the clustering scenarios, half of the subjects were singletons and the others shared a family member within the group (homogeneous) or across groups (heterogeneous). The distribution of $X_{1}$ was evenly distributed between 0 and 1. The distribution of $X_{2}$ was specified as normal with mean 0 or 1/2, respectively, with a standard deviation of 1, and an induced correlation between $X_{1}$ and $X_{2}$ of approximately 0.24. The residual standard error was also set to 1. Within each simulation, the null distribution was created using 2000 permutations.
Table LABEL:tbl-simresults displays the results from the simulation study. As expected, the models that do not appropriately account for clustering are anti-conservative (“lm homogeneous”, “naïve permutation homogeneous”) or overly conservative (“lm heterogeneous” and “naïve permutation heterogeneous”); the models are given in boldface. Methods that appropriately account for the clustering (e.g., “lme” or “correct permutation”) include the desired Type I error rate within the 95% CI. When there are cluster dependencies, the naïve method is not advised.
Table 1: Type I error rates for different scenarios (2000 simulations, each with 2000 permutations for each simulation). Models include MLR (lm), multiple linear mixed effects (lme), permutation not accounting for clustering (naïve), and permutation accounting for clustering (correct). Forms of clustering are independent (yellow), homogeneous (white), and heterogeneous (blue). The dichotomous treatment variable was permuted (Draper and Stoneman, 1996). Type I error was computed using $\alpha=0.05$.
Model/Scenario Type I error rate (95% CI) Results
lm independent 0.042 (0.034-0.052) includes desired alpha level
correct permutation independent 0.046 (0.037-0.056) includes desired alpha level
lm homogeneous 0.07 (0.059-0.082) anticonservative
lme homogeneous 0.051 (0.041-0.061) includes desired alpha level
naïve permutation homogeneous 0.072 (0.061-0.084) anticonservative
correct permutation homogeneous 0.045 (0.036-0.055) includes desired alpha level
lm heterogeneous 0.029 (0.023-0.038) conservative
lme heterogeneous 0.044 (0.035-0.054) includes desired alpha level
naïve permutation heterogeneous 0.034 (0.027-0.043) conservative
correct permutation heterogeneous 0.056 (0.047-0.068) includes desired alpha level
3 More ways to carry out Permutation Tests in MLR
The clustering setting demonstrates the importance of understanding exchangeability (and violations). However, we spend the rest of the paper describing exchangeability under the standard independent observation MLR model, because we believe that the discussion that follows is pedagogically quite powerful.
3.1 Why permutation tests?
As previously observed, even when both the normality conditions and the exchangeability are modestly violated, all four of the permutation test methods described below do reasonably well at controlling Type I errors (Anderson and Legendre 1999; Winkler et al. 2014). When error terms come from extremely skewed distributions (e.g., cubed exponentials) the permutation tests maintained the appropriate Type I error rate while normal theory methods were overly conservative (Anderson and Legendre 1999). There is strong evidence that in order to violate alpha-level testing, the data must violate the technical conditions to an extreme degree (Anderson and Legendre 1999; Tantawanich 2006; Winkler et al. 2014). Our own simulations (not shown) give the same results that despite differences across exchangeability conditions, the different permutation schemes (see Section 3.3) do not result in substantially violations of Type I errors.
If we anticipate that results will be indistinguishable, why is an understanding of permutation methods (and exchangeability) important? Why do we present the methods below as important for communicating exchangeability to our students?
  1. 1.
    We believe that exchangeability is a valuable concept for students to understand, as it undergirds key foundational knowledge of statistical inference.
  2. 2.
    There do exist settings where mathematical (i.e., based on the Central Limit Theorem) and permutation approaches produce different inferential results, even within the linear model framework. Knowledge of permutation-based approaches (plus computational skills) allows inference in such areas where linear model assumptions don’t hold, and permutation tests may be straightforward to implement in situations where typical parametric tests may exist but be extremely complicated. As long as permuted draws can be made from the appropriate null world, permutation tests can be used to make inferences without parametric assumptions.
3.2 Set-up
In order to work through the details of the permutation methods and corresponding exchangeability constraints, we provide notation and model specifications to describe the MLR model.
The population model of interest is
$E[Y|X_{1},X_{2}]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2},$
where $X_{1}$ is the variable of interest in predicting $Y$, and $X_{2}$ is a nuisance variable. The $\cdot$ (dot) notation indicates the other predictor variables included in the model.
The model estimated from the original dataset is given by
$\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
where $b_{1\cdot 2}$ is the sample coefficient on $X_{1}$ given that $X_{2}$ is in the model, and $b_{2\cdot 1}$ is the sample coefficient on $X_{2}$ given that $X_{1}$ is in the model.
An equivalent framing to the original data model is given for the permuted data. For example, if $Y$ is permuted to get $Y^{*}$, then the model estimated from the permuted dataset is given by
$\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
where $b^{*}_{1\cdot 2}$ is the sample coefficient on $X_{1}$ given that $X_{2}$ is in the model, and $b^{*}_{2\cdot 1}$ is the sample coefficient on $X_{2}$ given that $X_{1}$ is in the model, while it is $Y^{*}$ being regressed on $X_{1}$ and $X_{2}.$
All standard errors of the coefficients ($SE(b)$ and $SE(b^{*})$) are calculated using the normal ordinary least squares (OLS) formula used in standard linear regression software. Additionally, it is worth pointing out that underlying all of the methods (including OLS), the observations are assumed to be independent of one another.
3.3 Comparison of permutation methods
Unlike SLR, with MLR, there is no obvious choice of how to permute. In what follows, we discuss four different permutation methods and their exchangeability conditions. The goal of permuting is to create a null sampling distribution of the statistic of interest, here $b_{1\cdot 2},$ so that we may infer the variability of the statistic. That is, a distribution of the statistics under the setting where the outcome $Y$ and the predictor of interest $X_{1}$ are not linearly related. But also, the null distribution needs to be created under the exchangeability condition—that the permuting only leads to a change in the relationship between $Y$ and $X_{1}$ without changing any of the other variable relationships in the linear model. All four of the permuting schemes set the null hypothesis to $H_{0}\!\!:\beta_{1\cdot 2}=0.$
Statistics that are independent of all unknown parameters are called pivotal statistics. For permutation tests in the MLR setting, there are two reasons that we use t-statistics, which are pivotal or asymptotically pivotal (Winkler et al. 2014; ter Braak 1992). First, pivotal statistics allow for comparison across all methods, even the method that permutes the residuals of the full model and does not include $\beta_{1\cdot 2}=0$ in the test statistic. Second, the advantages of pivotal statistics are well-established (Hall and Titterington 1989; Hall and Wilson 1991; Westfall and Young 1993).
Using pivotal statistics, we dive into different permutation choices for the MLR model in order to help students develop a deeper understanding of the connection between exchangeability and how the permutation is implemented. Communicating the different permutation structures is an ideal way to explore the ideas of exchangeability. That is, even though the methods below are quite similar with respect to Type I errors and power, the discussion of the methods allows students study exchangeability and permutations more generally. We include extended details of the permutation schemes in Appendix 1. Here we briefly describe the four permutation methods with specific thought to the exchangeability violations for each one.
3.3.1 Permute $Y$
At first glance, it might seem like permuting the outcome variable would be a good way to break the relationship between $Y$ and $X_{1}$ (Manly 1986, 1997) (see complete algorithm in Section 1). Indeed, permuting $Y$ will break the relationship between $Y$ and $X_{1}$, which will force the null hypothesis to be true (which is what we want for testing). However, permuting $Y$ will also simultaneously break the relationship between $Y$ and $X_{2},$ which may not be acceptable if we need to preserve the relationship to mirror the original data structure. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $Y$. Note that when $Y$ is permuted, the original relationship between $X_{1}$ and $X_{2}$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{2}$ and $Y$, however, is problematic in terms of exchangeability. If we permute $Y,$ then exchangeability is met only if $Y$ and $X_{2}$ are uncorrelated in the original dataset.
Table 2: Different permutation schemes, variable relationships that are broken, and variable relationships that are preserved. Any relationships that are broken and not null violate exchangeability. Note that all methods require the observations to be independent of one another.
Permutation Broken Relationships Preserved Relationships Permutation distribution
Permute $Y$ $X_{1}$ $Y$ $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
Manly (1986), Manly (1997) $X_{2}$ $Y$ See Eq (1)
Permute $X_{1}$ $X_{1}$ $X_{2}$ $X_{2}$ $Y$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
N. Draper and Stoneman (1966) $X_{1}$ $Y$ See Eq (2)
Permute reduced model residuals $X_{1}$ $Y$ (if $X_{1}$ $X_{2}$ are uncorrelated) $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$
Freedman and Lane (1983) $X_{2}$ $Y$ See Eq (3)
Permute full model residuals None $X_{1}$ $X_{2}$ $t^{*}=\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$
ter Braak (1990) $X_{1}$ $Y$ See Eq (4)
ter Braak (1992) $X_{2}$ $Y$
3.3.2 Permute $X_{1}$
In order to maintain the relationship between $Y$ and $X_{2}$ (while still interested in the relationship between $Y$ and $X_{1}$), we might consider permuting $X_{1}$ instead of $Y$ (N. Draper and Stoneman 1966) (see complete algorithm in Section 1). Indeed, the permutation distribution created from permuting $X_{1}$ will force the null hypothesis to be true. However, permuting $X_{1}$ has the side effect that the relationship between $X_{1}$ and $X_{2}$ will be broken in the permuted data. If the data come from, for example, a randomized clinical trial (where $X_{1}$ is the treatment variable), then $X_{1}$ and $X_{2}$ will be independent in the original dataset, and permuting of $X_{1}$ will not violate the exchangeability condition. If $X_{1}$ and $X_{2}$ are correlated in the original dataset, as seen in our AP course and socio-economic status example, then permuting $X_{1}$ violates the exchangeability condition. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $X_{1}$. Note that when $X_{1}$ is permuted, the original relationship between $X_{1}$ and $Y$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $X_{2}$, however, is problematic under general conditions.
3.3.3 Permute reduced model residuals
In order to address some of the concerns of the first two MLR permutation methods, Freedman and Lane (1983) propose a permutation scheme based on residuals (see complete algorithm in Section 1). The permutation preserves the relationship between $X_{1}$ $X_{2}$ as well as the relationship between $X_{2}$ $Y$. However, in order for the relationship between $X_{1}$ $Y$ to be broken (i.e., to obtain a null sampling distribution for the test of $H_{0}:\beta_{1\cdot 2}=0),$ $X_{1}$ and $X_{2}$ must not be associated. Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that when permuting the reduced model residuals, the original relationships between both $X_{1}$ $X_{2}$ and additionally between $X_{2}$ $Y$ are preserved (which is what we want in terms of exchangeability). The null hypothesis is true only if the relationship between $X_{1}$ and $Y$ is broken, and that happens only when $X_{1}$ and $X_{2}$ are uncorrelated. Appendix Section 2 sketches the dependence of $\rho(Y^{*},X_{1})$ (the correlation between $Y^{*}$ and $X_{1}$) on $\rho(X_{1},X_{2})$ (the correlation between $X_{1}$ and $X_{2}).$
3.3.4 Permute full model residuals
As an extension to Freedman and Lane (1983), ter Braak’s (1990, 1992) permutation method permutes the residuals from the full model (see complete algorithm in Section 1). Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the full model residuals. Note that permuting the residuals under the full model allows all of the exchangeability conditions to hold. The new model does not force the null hypothesis to be true, which is why the test statistic measures the deviation of the permuted coefficients (on $X_{1}$), $b_{1\cdot 2}$, to the original data model coefficient (on $X_{1}$), $b^{*}_{1\cdot 2}$, instead of comparing to zero.
4 Discussion and pedagogical recommendations
In the specific context of the MLR model, there are violations of exchangeability (e.g., clustering) which affect performance and violations of exchangeability (e.g., the correlation between $X_{1}$ and $X_{2}$) which do not affect performance. Our deep dive into permutation tests for MLR is meant to communicate ideas of exchangeability mathematically and pedagogically. We acknowledge that we do not have a smoking gun example which shows which of the standard MLR permutation methods is “best” (indeed, they are roughly equivalent procedures), but we find their introduction to students as a helpful structure to explore exchangeability in a meaningful way. However, an advantage of permutation tests is that they give accurate inferences even in small samples (Anderson and Legendre 1999).
Our work presents some of the existing literature on permutation tests in the MLR setting. While it is not immediately obvious how to best permute an MLR model, it turns out that, generally, the different methods perform similarly with respect to Type I errors (and power). However, the difficult and extremely powerful concept of exchangeability can be accessed through the MLR setting in the classroom. After digging into the exchangeability ideas for the MLR case, students are able to apply a permutation approach to more complicated data settings (which are often encountered in real applications), like clustered observations. This guidance is consistent with the advice of Chance et al. (2024), who hoped “to help instructors understand the differences in the simulation strategies to better inform their own decisions of how to adapt a simulation approach for their classes and to better respond to student questions that may arise”.
Another compelling rationale to devote time to these questions is the concept that “the test follows from the design”:
There is a clear logical link between the statistical test we use and the experimental design we opted for: using a permutation test is entirely warranted by the random assignment of participants to two equal-sized groups. Stressing the link between experimental design and statistical inference – rather than considering them separately – is of huge pedagogical, as well as practical, use, I believe (Vanhove 2015).
We close with some practical guidance about how to bring exchangeability and permutation tests into the classroom. The suggestions are organized by what type of classroom might be most appropriate.
introductory undergraduate
  • Teacher creates a simulation program (Rmd/qmd file or Shiny App) where students can change the error structure within an MLR analysis and discover the desired Type I error rates are not achieved only in the case when the errors are particularly egregious.
intermediate / advanced undergraduate
  • Teach permutation tests for MLR. Talk to your students about the choice they will have to make (permute $Y$? $X_{1}$? residuals?) and that all modeling contains choices. Choices are not usually objective or unbiased, so an ability to defend their choice is what gives them power as a statistician.
  • Describe the difference between creating a null sampling distribution (i.e., making sure that the null hypothesis is true) and establishing that exchangeability holds (i.e., making sure that the Type I error rates will be accurate). Have students describe which aspect of the permutation does which job.
  • Come up with scenarios (or use our cluster scenarios!) where the students can figure out the correct permutation schemes. See an example from JH’s class that uses permutations to address a research question using a stratified two-sample test. https://st47s.com/Math154/Notes/permschp.html#macnell-teaching-evaluations-stratified-two-sample-t-test
  • Carry out a variety of permutation tests using a multiple linear regression model from an example dataset (e.g., the bridges data in the supplementary materials). Have students check the implementations or provide some of the implementations and have them carry out the others.
advanced undergraduate
  • Ask your students what it means that “the test follows from the design.” Use the cluster scenarios and apply the opposite permutation scheme. Have them compare the permutation scheme and the experimental design to figure out what makes most sense in terms of the relationship of the permuting and the design.
  • Ask students to read the paper by Helwig (2019a) and explore the implications for the MLR setting via the nptest package in R (Helwig 2023). Have them explore how and why robust permutation tests (DiCiccio and Romano 2017) use alternative test statistics.
advanced undergraduate / beginning graduate
  • Present some or all of the MLR methods described here (with more in Winkler et al. (2014)) and have students fill out a blank Table LABEL:tbl-perms. Ask students to report which relationships are broken and which are preserved.
  • Find permutation tests in the literature and have students describe how the permutation scheme upholds (or doesn’t!) the exchangeability condition.
graduate
  • Find theoretical work that proves exchangeability and demonstrate the important mapping of the theory to the applied problems (or maybe the exchangeability is not vital to the performance of the method?).
We suggest some ideas at particular levels, but many of the pedagogical ideas can be adjusted to be effective at a variety of levels. Whether to permute $Y$ or $X_{1}$ can be taught as early as introductory statistics, where students have done some hypothesis testing and some multivariate modeling. Graduate students can think carefully about how to prove that conditions of exchangeability are met. The group of students who might most benefit from the ideas we’ve presented include upper level undergraduates and early graduate students who are focused on understanding which models are best in which settings and how to differentiate those settings.
5 Acknowledgements
This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. We thank Shiya Cao and Lindsay Poirier for pointing us toward the bridges data and Shiya Cao for useful comments on an earlier draft of the manuscript and the reviewers and associate editor for helpful suggestions.
6 Disclosure statement
The authors report there are no competing interests to declare.
7 Reproducibility statement
The code used to undertake the simulation study reported in Table LABEL:tbl-simresults and the supplementary resources (bridges analysis) has been made available at https://osf.io/7zcfu/?view.
Appendix 0 Appendix
Here we review different approaches for carrying out permutation tests for multiple linear regression models.
1 Permutation algorithms
Permute $Y$
Summary of algorithm when permuting $Y$ (Manly)
  1. 1.
    Fit the original model and obtain coefficient estimates ($b_{0\cdot 1,2}$, $b_{1\cdot 2}$, and $b_{2\cdot 1}$) and corresponding standard error estimates ($SE(b_{0\cdot 1,2})$, $SE(b_{1\cdot 2})$, and $SE(b_{2\cdot 1})$):
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Permute $Y$ to obtain $Y^{*}$.
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates ($b^{*}_{0\cdot 1,2}$, $b^{*}_{1\cdot 2}$, and $b^{*}_{2\cdot 1}$) and corresponding standard error estimates ($SE(b^{*}_{0\cdot 1,2})$, $SE(b^{*}_{1\cdot 2})$, and $SE(b^{*}_{2\cdot 1})$):
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times. For example, $P$ = 1000.
  5. 5.
    From the $P$ copies of $b^{*}_{1\cdot 2}$ and $P$ copies of $SE(b^{*}_{1\cdot 2})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (1)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
Permute $X_{1}$
Summary of algorithm when permuting $X_{1}$ (Draper and Stoneman)
  1. 1.
    Fit the original model and obtain coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Permute $X_{1}$ to obtain $X^{*}_{1}$.
  3. 3.
    Fit a model on the permuted $X_{1}^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X^{*}_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (2)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
Permute reduced model residuals
Summary of algorithm when permuting reduced model residuals (Freedman and Lane)
  1. 1.
    Fit the original model on $X_{2}$ only and obtain coefficient estimates and corresponding standard error estimates of the reduced model:
    $\widehat{Y}=b_{0\cdot 2}+b_{2}X_{2}$
  2. 2.
    Let the residuals $R_{Y\cdot 2}=Y-b_{0\cdot 2}-b_{2}X_{2}$, and permute $R_{Y\cdot 2}$ to obtain $R^{*}_{Y\cdot 2}.$ Define the permuted outcome variable as $Y^{*}=b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2}.$
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (3)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
It is not immediately obvious that a high correlation between $X_{1}$ and $X_{2}$ leads to a non-null distribution. We provide evidence for the claim in two ways: by describing the permutation scheme and by giving a sketch for the proof of the claim.
Claim. The correlation between $Y^{*}$ and $X_{1}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
  1. (1)
    The permutation scheme is designed to create new values for the coefficient on $X_{1}$ under the setting that $X_{1}$ and $Y$ are uncorrelated given $X_{2}$. Repeated sets of the “permuted” $Y^{*}$ values are generated by adding noise (permuted residuals) to the fitted values from the model on $X_{2}$ only. At first glance, it seems as though the method creates $Y^{*}$ values that are associated with $X_{2}$ (because they built from the $X_{2}$-model fitted values) and not associated with $X_{1}$ (because they are built from a model that completely ignores $X_{1}$). However, if $X_{1}$ and $X_{2}$ are correlated, then $Y^{*}$ values that are correlated with $X_{2}$ will naturally be correlated with $X_{1}$.
  2. (2)
    In Appendix 2, we give a proof sketch outlining the derivation for the covariance between $Y^{*}$ and $X_{1}$. Although the proof is only outlined, the dependence on the correlation between $X_{1}$ and $X_{2}$ is clear. The approximate value of $\rho(Y^{*},X_{1})$ derived in Appendix 2 is validated empirically (results not shown); see Ye (2023) for empirical results.
Permute full model residuals
Summary of algorithm when permuting full model residuals (ter Braak)
  1. 1.
    Fit the original model and obtain coefficient estimates and corresponding standard error estimates of the full model:
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Let the residuals $R_{Y\cdot 1,2}=Y-b_{0\cdot 1,2}-b_{1\cdot 2}X_{1}-b_{2\cdot 1}X_{2}$, and permute $R_{Y\cdot 1,2}$ to obtain $R^{*}_{Y\cdot 1,2}.$ Define the permuted outcome variable as $Y^{*}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}+R^{*}_{Y\cdot 1,2}.$
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$ (4)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
The full model residual method has a bootstrap flavor, but the permutation of the residuals is done without replacement (whereas bootstrapping is done with replacement). Similar to the development of other bootstrapping methods (Efron and Tibshirani 1994, 87), the underlying technical condition for ter Braak’s method is that:
$F(x)_{\big{(}\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}\big{)}}\approx F(x)_{\big{(}\frac{b_{1\cdot 2}-\beta_{1\cdot 2}}{SE(b_{1\cdot 2})}\big{)}},$ (5)
where $F$ is the cumulative distribution function of the random variable described in the subscript of the function.
The sampling distribution of $t^{*}=\frac{b^{*}_{1\cdot 2}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$ approximates the sampling distribution of $t=\frac{b_{1\cdot 2}-\beta_{1\cdot 2}}{SE(b_{1\cdot 2})}$ due to the theoretical underpinnings from bootstrapping. That is, the variability of $b_{1\cdot 2}^{*}$ around $b_{1\cdot 2}$ mimics the variability of $b_{1\cdot 2}$ around $\beta_{1\cdot 2}.$
Under Equation (5), we can use the $t^{*}$ distribution constructed from many permutations of the same dataset to carry out a hypothesis test. If $H_{0}\!\!:\beta_{1\cdot 2}=0$ is true, then $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$ would be a likely value in the $t^{*}$ distribution, corresponding to a non-significant $p$-value, resulting in a failure to reject $H_{0}.$
However, if $H_{A}\!\!:\beta_{1\cdot 2}\neq 0$ is true and, say, $\beta_{1\cdot 2}=47,$ then we would expect $\hat{t}=\frac{b_{1\cdot 2}-47}{SE(b_{1\cdot 2})}$ to lie well within the $t^{*}$ distribution, while $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$ would lie on the margins, leading to a small $p$-value that concludes the test by rejecting $H_{0}.$
Furthermore, ter Braak (1990) synthesizes equations from various authors – Efron (1982) for the bootstrap, along with Cox and Hinkley (1974) and Lehmann and D’Abrera (1975) for the permutation – to make the following statements regarding the expected value and variance of the estimated slope coefficients in the bootstrapping versus permutation settings. $b_{1\cdot 2}^{+}$ corresponds to the coefficient derived from a standard bootstrap (full model), while $b_{1\cdot 2}^{*}$ corresponds to the permutation of the residuals under the full model.
$E^{+}(b_{1\cdot 2}^{+})$ $=E^{*}(b_{1\cdot 2}^{*})=b_{1\cdot 2}$ (6)
$var^{+}(b_{1\cdot 2}^{+})$ $=(1-1/n)var^{*}(b_{1\cdot 2}^{*}).$ (7)
Equation (6) indicates that the expected values of $b_{1\cdot 2}^{+}$ and $b_{1\cdot 2}^{*}$ are both $b_{1\cdot 2}$. Equation (7) suggests that the variance of $b_{1\cdot 2}^{+},$ the bootstrapped estimate, is smaller than the variance of $b_{1\cdot 2}^{*}.$ Hence, ter Braak uses Equations (6) and (7), along with the order property that $b_{1\cdot 2}^{+}$ and $b_{1\cdot 2}^{*}$ differ by $O(1/n)$ in second or higher order moments, to justify his proposal of the full model residual permutation strategy.
2 Impact of correlation between $X_{1}$ and $X_{2}$
Claim. The correlation between $Y^{*}$ and $X_{1}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
Proof sketch.
Throughout the proof sketch, there are places where we have simplified the argument by considering statistics to be fixed parameters (not an unreasonable approximation under large sample sizes where Slutsky’s Theorem holds (Slutsky 1925)).
$\rho(Y^{*},X_{1})$ $=\frac{cov(Y^{*},X_{1})}{\sqrt{var(Y^{*})var(X_{1})}}$
Breaking down each part of the correlation between $Y^{*}$ and $X_{1}$, we compute both $cov(Y^{*},X_{1})$ and $var(Y^{*}).$
$cov(Y^{*},X_{1})$ $=cov(b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2},X_{1})$
$=cov(b_{0\cdot 2},X_{1})+cov(b_{2}X_{2},X_{1})+cov(R^{*}_{Y\cdot 2},X_{1})$
$\approx cov(b_{2}X_{2},X_{1})$ (8)
$\approx b_{2}\cdot cov(X_{1},X_{2})$ (9)
$=b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}$
Note that $b_{0\cdot 2}$ and $b_{2}$ are random variables because they are statistics, so Equations (9,10,11) are all approximate. In Equation (8), we assume that $b_{0\cdot 2}$ and $R^{*}_{Y\cdot 2}$ are both independent of $X_{1}$, which leads to $cov(b_{0\cdot 2},X_{1})=cov(R^{*}_{Y\cdot 2},X_{1})=0$.
$var(Y^{*})$ $=var(b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2})$
$\approx(b_{2})^{2}var(X_{2})+var(R^{*}_{Y\cdot 2})$ (10)
$var(R^{*}_{Y\cdot 2})$ is broken down into pieces.
$var(R^{*}_{Y\cdot 2})$ $=var(R_{Y\cdot 2})$
$=var(Y-b_{0\cdot 2}-b_{2}X_{2})$
$\approx var(Y)+(b_{2})^{2}var(X_{2})-2b_{2}cov(Y,X_{2})$ (11)
$cov(Y,X_{2})$ $=cov(\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2}+\varepsilon,X_{2})$
[…]
$=\beta_{2\cdot 1}var(X_{2})+\beta_{1\cdot 2}cov(X_{1},X_{2})$ (12)
Plugging $cov(Y,X_{2})$ into $var(R^{*}_{Y\cdot 2})$ and $var(R^{*}_{Y\cdot 2})$ into $var(Y^{*})$, we can approximate $var(Y^{*}).$
[…]
$=2b_{2}(b_{2}-\beta_{2\cdot 1})var(X_{2})+var(Y)-2b_{2}\beta_{1\cdot 2}cov(X_{1},X_{2})$
$=2b_{2}[(b_{2}-\beta_{2\cdot 1})var(X_{2})-\beta_{1\cdot 2}cov(X_{1},X_{2})]+var(Y)$
Putting it all together gives:
$\rho(Y^{*},X_{1})$ $=\frac{cov(Y^{*},X_{1})}{\sqrt{var(Y^{*})var(X_{1})}}$
$\approx\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}}{\sqrt{var(Y^{*})var(X_{1})}}$
$=\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{2})}}{\sqrt{var(Y^{*})}}$
$=\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{2})}}{\sqrt{2b_{2}[(b_{2}-\beta_{2\cdot 1})var(X_{2})-\beta_{1\cdot 2}cov(X_{1},X_{2})]+var(Y)}}.$ (13)